### v1

# Flexible game-playing AI with AlphaViT: adapting to multiple games and board sizes

Abstract
This paper presents novel game AI agents based on the AlphaZero framework, enhanced with Vision Transformers (ViT): AlphaViT, AlphaViD, and AlphaVDA. These agents are designed to play various board games of different sizes using a single model, overcoming AlphaZero’s limitation of being restricted to a fixed board size. AlphaViT uses only a transformer encoder, while AlphaViD and AlphaVDA contain both an encoder and a decoder. AlphaViD’s decoder receives input from the encoder output, while AlphaVDA uses a learnable matrix as decoder input. Using the AlphaZero framework, the three proposed methods demonstrate their versatility in different game environments, including Connect4, Gomoku, and Othello. Experimental results show that these agents, whether trained on a single game or on multiple games simultaneously, consistently outperform traditional algorithms such as Minimax and Monte Carlo tree search using a single DNN with shared weights, while approaching the performance of AlphaZero. In particular, AlphaViT and AlphaViD show strong performance across games, with AlphaViD benefiting from an additional decoder layer that enhances its ability to adapt to different action spaces and board sizes. These results may suggest the potential of transformer-based architectures to develop more flexible and robust game AI agents capable of excelling in multiple games and dynamic environments.
1 Introduction
In recent years, artificial intelligence (AI) has made remarkable progress, demonstrating its potential in a wide range of applications. One application area where AI has shown significant prowess is in mastering board games, surpassing the skills of top human players in many games. Historical achievements include AI outperforming humans in games such as checkers, chess (Campbell et al., 2002), and Othello (Buro, 1997). A significant milestone was reached in 2016 when AlphaGo (Silver et al., 2016), an AI specialized in the game of Go, defeated one of the world’s top players. The subsequent introduction of AlphaZero (Silver et al., 2018), capable of mastering multiple board games such as Chess, Shogi, and Go, has further solidified the superhuman capabilities of AI in this domain.
However, current game-playing AI agents have a significant drawback: they are designed to play only one specific game and cannot play other games. Even if the rules remain the same, they cannot handle variations in board size. In contrast, humans can easily switch between different board sizes. For instance, Go beginners often start by practicing on smaller boards (e.g., $9\times 9$) before moving on to larger boards (e.g., $19\times 19$). However, AI agents such as AlphaZero, which are designed for a specific game and fixed board size, cannot adapt to these changes without significant reprogramming.
For AlphaZero specifically, the core of this limitation lies in the architecture of AlphaZero’s deep neural network (DNN), which requires a fixed input size. AlphaZero uses a DNN consisting of residual blocks and multilayer perceptrons (MLPs). These components are designed for a fixed input size. The output size of the residual blocks varies as the input size changes, creating an inconsistency with the expected output size for MLPs. As a result, AlphaZero cannot function properly when faced with even small variations in board size. To address this issue, we propose to replace the residual blocks in the AlphaZero framework with Vision Transformers (ViT) (Dosovitskiy et al., 2021).
ViT is an image classification DNN based on the transformer architecture. ViT divides an image into several patches and infers a class of images from these patches. A key advantage of ViT is its ability to process images independently of input size, allowing for flexible adaptation to different board sizes within the AlphaZero framework.
In this study, we present game-playing agents that use a single DNN to handle multiple games and variable board sizes. These agents, named AlphaViT, AlphaViD (AlphaViT with a transformer decoder layer), and AlphaVDA (AlphaViD with learnable action tokens), are based on the AlphaZero framework. The agents predict the value and policy of game states using a DNN, while decisions are made using Monte Carlo Tree Search (MCTS). Our computational experiments show that these models can simultaneously play games such as Connect4, Gomoku, and Othello using a single DNN with shared weights. Moreover, the proposed agents outperform traditional algorithms such as Minimax and MCTS in various games, while approaching the performance of AlphaZero, whether trained on a single game or on multiple games simultaneously.
2 Related work
Game-playing AI agents have achieved superhuman-level performance in traditional board games such as Checkers (Schaeffer et al., 1993), Othello (Buro, 1997, 2003), and Chess (Campbell, 1999; Hsu, 1999; Campbell et al., 2002). In 2016, AlphaGo (Silver et al., 2016), a Go-playing AI, has defeated the world’s top Go player, becoming the first superhuman-level Go-playing AI. AlphaGo relies on supervised learning from a large database of expert human moves and self-play data. Subsequently, AlphaGo Zero (Silver et al., 2017) has defeated AlphaGo without preparing a large training dataset. In 2018, Silver et al. (2018) have proposed AlphaZero, which has no restrictions on playable games. AlphaZero has outperformed other superhuman-level AIs at Go, Shogi, and Chess. Interestingly, AlphaZero’s versatility extends beyond traditional two-player perfect information games. Research has explored its potential in more complex scenarios. For example, Hsueh et al. (2018) have shown AlphaZero’s potential in nondeterministic games. Other extensions include handling continuous action spaces (Moerland et al., 2018) and support for multiplayer games (Petosa and Balch, 2019). However, AlphaZero cannot play various games simultaneously or handle games with the same rules but different board sizes using a single DNN with shared weights.
This limitation is due to the use of the DNN in AlphaZero for policy and value estimation. AlphaZero’s DNN consists of residual blocks and MLPs. While the residual blocks excel at extracting features from input images, the MLPs are constrained by a fixed input size. As a result, the DNN is not sufficiently flexible to accommodate variations in board size. To address this limitation, the integration of Vision Transformer (ViT) (Dosovitskiy et al., 2021) into AlphaZero may provide a solution.
The transformer architecture, initially designed for natural language processing (Vaswani et al., 2017), has shown remarkable effectiveness in various domains. The transformer architecture has also been successfully applied to image-processing tasks. Transformer-based models have achieved exceptional performance in various image-related tasks, including image classification (Dosovitskiy et al., 2021), semantic segmentation (Xie et al., 2024), video classification (Li et al., 2022), and video captioning (Zhao et al., 2022). ViT, introduced by Dosovitskiy et al. (Dosovitskiy et al., 2021), is a remarkable example of a transformer-based model for image processing. ViT has achieved state-of-the-art performance in image classification at the time of its introduction.
A key feature of ViT is its independence from the size of the input image (Dosovitskiy et al., 2021). Unlike convolutional neural networks (CNNs), which require fixed-size inputs, ViT can handle images of various sizes by dividing them into fixed-size patches and treating each patch as a token in the transformer architecture. This flexibility allows ViT to be highly adaptable and efficient in handling different image sizes.
While AlphaZero can only play the game it was trained on, humans can play multiple games with a single brain, such as Chess, Othello, and Connect4. In addition, humans can adapt to different board sizes if they know the rules of a game. However, even if only the board size of a game changes, AI such as AlphaZero cannot play the game. The goal of this study is to overcome this limitation by developing AI agents based on AlphaZero, which can be generalized across different board sizes and various games.
3 AlphaViT, AlphaViD, and AlphaVDA
AlphaZero’s game-playing capability is constrained to games with identical board sizes and rules as those used during its training. This limitation is due to AlphaZero’s ResNet-based deep neural network (DNN), which consists of residual blocks followed by multilayer perceptrons (MLPs) for value and policy computation. The MLPs receive input from the final residual block, assuming a fixed input size. Consequently, AlphaZero’s DNN cannot adapt to variations in board size. Therefore, AlphaZero’s performance is limited to games with fixed board size, restricting its performance to games with predetermined dimensions.
To overcome this limitation, AlphaViT, AlphaViD, and AlphaVDA have been developed as game-playing AIs based on AlphaZero but using Vision Transformer (ViT) architecture. These game-playing AI agents use a combination of a DNN and MCTS (Fig. 1). The DNN receives board states and outputs value estimates and move probabilities (policies). The MCTS then searches the game tree using these outputs. The MCTS searches a game tree using the estimated value and move probability. By incorporating ViT instead of residual blocks, AlphaViT, AlphaViD, and AlphaVDA can overcome the limitation of AlphaZero and can play games with different board sizes and rules.
An overview of AlphaViT’s DNN architecture is shown in Fig. 2. The DNN of AlphaViT is based on ViT, which has no image-size limitation and can classify an image even if the input image size differs from the training image size. This flexibility allows AlphaViT and AlphaViD to play games with different board sizes using the same network.
In AlphaViT, the boards are input to ViT. Initially, these inputs are transformed into patch embeddings using a convolutional layer. Using the convolution layer allows for easy adjustment of patch division parameters such as patch size, stride, and padding. The final outputs of the encoder layer are then used to compute value and policy estimates.
AlphaViD and AlphaVDA incorporate both transformer encoder and decoder layers for value and move probability calculations. In AlphaViD, the decoder layer receives input derived from the encoder layer’s output. Conversely, AlphaVDA utilizes learnable embeddings as input for its decoder layer.
Importantly, AlphaViT, AlphaViD, and AlphaVDA can play any game that AlphaZero can, as they employ the same fundamental game-playing algorithm. Their enhanced flexibility in handling various board sizes and game rules represents a significant advancement in AI game-playing capabilities.
Figure 1: AlphaViT and AlphaViD use the same procedure as AlphaZero. They receive board states and a current player, and decide a move using MCTS. MCTS searches a game tree using the value and policy estimated by the DNN.
Figure 2: Overview of the AlphaViT model architecture. A board state is divided into fixed-size patches and linearly embedded by a convolutional layer. Position embeddings are added to these patch embeddings. The resulting sequence of patch embeddings is appended with the value embedding and the game embedding. The sequence is fed to a transformer encoder. Finally, the value and the move probability are estimated using a multilayer perceptron (MLP) that takes the outputs from the last layer of the transformer encoder as input.
3.1 Architectures of DNNs
AlphaViT
AlphaViT’s deep neural network (DNN) predicts the value $v(s)$ and the move probability distribution $\bm{p}$ with components $p(a|s)$ for each action $a$, given a game state $s$. In a board game, $s$ represents the state of a board, and $a$ denotes a move.
The input to the DNN is an $H\times W\times(2T+1)$ image stack ${\rm\bm{x}}\in R^{H\times W\times(2T+1)}$, consisting consisting of $12T+$ 1 binary feature planes of size $H\times W$. Here, $H$ and $W$ are the dimensions of the board, and $T$ is the number of history planes. The first $T$ feature planes represent the occupancies of the first player’s discs, where a feature value of 1 indicates that a disc occupies the corresponding cell and 0 otherwise. The following $T$ feature planes represent the occupancies of the second player’s discs. The final feature level represents the disc color of the current player, where 1 and -1 represent the first and second players, respectively.
The convolutional layer processes the image stack ${\rm\bm{x}}$ to generate the patch embeddings $x_{p}$. This layer divides the image stack into $P\times P$ image patches with stride $s$ and padding $p$ and reshapes them into flattened 2D patches ${\rm\bm{x}}_{p}\in R^{(WH)\times N_{e}}$, where $N_{e}$ is the embedding size. These flattened patches are treated as a sequence of token embeddings ${\rm\bm{z}}_{0}$:
${\rm\bm{z}}_{0}={\rm\bm{x}}=[{\rm\bm{x}}^{1}_{p};...;\bm{x}^{i}_{p};...;\bm{x}^{N}_{p}],$ (1)
where ${\rm\bm{x}}^{i}_{p}$ is the $i$th token embedding in the sequence.
To preserve position information, learnable 2D position embeddings $E_{\mathrm{pos}}$ are added to the patch embeddings. The size of these positional embeddings is fixed and scaled based on the board size and hyperparameters to ensure compatibility with patch embeddings ${\rm\bm{x}}_{0}$. The position embeddings are incorporated into the token embeddings ${\rm\bm{x}}_{0}$ as follows:
${\rm\bm{z}}_{0}\leftarrow{\rm\bm{z}}_{0}+E_{\mathrm{pos}}.$ (2)
The output size of a transformer encoder layer corresponds to the number of input embeddings. For Gomoku, where the action space is $HW$, AlphaViT requires $HW$ embeddings. To achieve this, we set $k=2n+1$, where $n$ is a positive integer, stride $s=1$, and padding $pad=\lfloor k/2\rfloor$.
For Othello, the action space is $HW+1$ to accommodate the pass move. Using the same parameters as Gomoku ($k=2n+1$, $s=1$, and $pad=\lfloor k/2\rfloor$), AlphaViT needs one additional embedding for the pass move. To address this, we introduce a learnable pass token $\bm{x}_{\mathrm{pass}}$ and append it to the input embeddings.
To estimate the board value, we prepend a learnable value embedding $\bm{x}_{\mathrm{value}}$. Additionally, to enable AlphaViT to distinguish between different game types, we incorporate a static game embedding $\bm{x}_{\mathrm{game}}$, represented using one-hot encoding. These embeddings are appended to the initial embeddings ${\rm\bm{z}}_{0}$, resulting in
${\rm\bm{z}}_{0}=[\bm{x}_{\mathrm{value}};\bm{x}_{\mathrm{game}};\bm{x}^{1}_{p}\bm{\mathrm{E}};...;\bm{x}^{WH}_{p}\bm{\mathrm{E}};\bm{x}_{\mathrm{pass}}].$ (3)
The position embeddings are added before the pass, value, and game embeddings are appended. This approach allows for scaling of the position embeddings without affecting these embeddings, which do not contain positional information. The resulting sequence ${\rm\bm{z}}_{0}$ serves as the input for the transformer encoder.
Sequence $\bm{z}_{0}$ is fed into the transformer encoder, which consists of $L$ transformer encoder layers. The output of the last encoder layer ${\rm\bm{z}}_{L}$ consists of $M\times N+2$ vectors. The first vector $z^{0}_{L}$ derived from the value embedding is processed by the value head implemented as a multilayer perceptron (MLP) denoted as $\text{MLP}_{v}$. This head estimates the value $v$:
$v=\tanh(\mathrm{NLP}_{v}(LN(z^{0}_{L}))),$ (4)
where LN denotes layer normalization. Tanh activation ensures the value ranges from $-1$ to $1$, representing the winning probability.
The vectors $z^{2}_{L},...,x^{WH+1}_{L}$ derived from the board patches are processed by the policy head, another MLP ($\mathrm{NLP}_{p}$). The policy head outputs the move probability $p_{m,n}$ for each board position $(m,n),0\leq m<H,0\leq n<W$ using the sigmoid function. The policy head outputs matrix $R^{M\times N}$ representing the move probabilities:
$p(m,n)=Sigmoid(\mathrm{MLP}_{p(m,n)}(LN(z^{m*H+n+1}_{L}))).$ (5)
For Othello, $p(m,n-1)$ represents the probability of pass action. For Connect4, since a player can select only $W$ actions, AlphaViT uses only $p(0,n)$.
To decide on a move, AlphaViT employs Monte Carlo Tree Search with Upper Confidence Bound (UCT), using the value and move probability predicted by the DNN. AlphaViT uses the same MCTS algorithm as AlphaZero, as shown in Appendix 1. The parameters used in AlphaViT are presented in Appendix 3.
AlphaViD and AlphaVDA
AlphaViT has a significant drawback: the size of the policy vector is fixed to the input size of the transformer encoder layer. To address this problem, we propose an improved AlphaViT called AlphaViD. Although similar to AlphaViT in many respects, AlphaViD has a transformer decoder, as shown on the left in Fig. 3. AlphaViD uses the transformer encoder and decoder to estimate the value and move probability, respectively.
The input board is linearly embedded using a convolutional layer and fed into a transformer encoder, similar to AlphaViT. The input embedding sequence is as follows
${\rm\bm{z}}_{0}=[\bm{x}_{\mathrm{value}};\bm{x}_{\mathrm{game}};\bm{x}^{1}_{p}\bm{\mathrm{E}};...;\bm{x}^{WH}_{p}\bm{\mathrm{E}}].$ (6)
The embedding composition of AlphaViD is different from that of AlphaViT. Since AlphaViD only estimates the value from the encoder output, it does not require the embedding size to exceed the action space size. The embedding sequence consists of value and game tokens and patch embeddings. The estimated value is obtained using the value head, which processes the output corresponding to the value embedding from the last layer of the transformer encoder.
In AlphaViD, the move probability is estimated using the transformer decoder and MLPp. The input embeddings for the transformer decoder are generated from the outputs of the transformer encoder corresponding to the patch embeddings processed through a fully connected layer:
$E^{\prime}_{d}=MLP([\bm{z}^{2}_{L_{e}};...;\bm{z}^{n*H+m+1}_{L_{e}}]),E^{\prime}_{d}\in R^{HW\times N_{e}}.$ (7)
Since the sequence size must match the input size of the transformer decoder, $E^{\prime}_{d}$ is interpolated to $E_{d}\in\mathbb{R}^{N_{a}\times N_{e}}$, where $N_{a}$ is the action space size and $N_{e}$ is the embedding size of the transformer decoder. This allows flexibility in adjusting the action size depending on the game type and board size. The transformer decoder receives $E_{d}$ and the output of the transformer encoder $z^{L}$, just like the original transformer. MLPp calculates the move probability from the output of the last layer of the transformer decoder $y_{L}$.
AlphaVDA has the architecture shown on the right in Fig. 3. AlphaVDA has a similar architecture to AlphaViD, but uses learnable embeddings $E^{\prime}_{d}$ as input to the transformer decoder. These input embeddings are interpolated to fit the action size.
Figure 3: Architectures of the AlphaViD and AlphaVDA models. The input board state is divided into fixed-size patches and linearly embedded using a convolutional layer. Position embeddings are added to the resulting patch embeddings to preserve spatial information. The sequence of patch embeddings is then concatenated with a learnable value embedding and a static game embedding. This input sequence is fed into a transformer encoder. The estimated value is obtained through a multilayer perceptron (MLP) that takes the output corresponding to the value embedding from the last layer of the transformer encoder. The outputs corresponding to the patch embeddings are passed through a fully connected layer to generate input embeddings for the transformer decoder. Finally, the move probabilities are produced by another MLP that takes the output of the last layer of the transformer decoder. For AlphaVDM, the input embeddings of the transformer decoder is learnable embeddings.
3.2 AlphaViT, AlphaViD, and AlphaVDA with ResNet
AlphaViT, AlphaViD, and AlphaVDA use linear transformation to make tokens for transformer encoder. The tokenizer is not limited to linear transformation. In this study, we evaluate ResNet instead of linear transformation when it is used for tokenizer.
4 Experimental settings
4.1 Games
This study evaluates the performance of AlphaViT and AlphaViD in six games: Connect4, Connect4 5x4, Gomoku, Gomoku 6x6, Othello, and Othello 6x6. These games are two-player, deterministic, zero-sum games with perfect information. Connect4, published by Milton, is a connection game played on a $7\times 6$ board. Players take turns dropping discs onto the board. A player wins by forming a straight line of four discs horizontally, vertically, or diagonally. 54Connect4 is Connect4 with a $5\times 4$ board. Gomoku is a connection game in which players place stones on a board to form a line of five stones in a row, either horizontally, vertically, or diagonally. This study uses a $9\times 9$ board for Gomoku and a $6\times 6$ board for Gomoku 6x6. Othello (Reversi) is a two-player strategy game played on an $8\times 8$ board. In Othello, the disc is white on one side and black on the other. Players take turns placing a disc with their assigned color facing up. During a game, discs of the opponent’s color are flipped to the current player’s color if they are in a straight line and bounded by the disc just placed and another disc of the current player’s color. 66Othello is played on a $6\times 6$ board in this study.
4.2 Opponents
This study evaluates the performance of AlphaViT and AlphaViD using five different AI methods: AlphaZero, two variants of Monte Carlo Tree Search (MCTS) labeled MCTS100 and MCTS400, Minimax, and Random. AlphaZero is trained using the method described in Appendix 1. The MCTS methods (MCTS100 and MCTS400) were implemented using different numbers of simulations (100 and 400, respectively). The details of MCTS can be found in Appendix LABEL:sec:mcts. In these MCTS methods, the child nodes are expanded at the fifth visit to a node. Minimax selects a move by the minimax algorithm based on the evaluation table described in Appendix 4. Random selects moves uniformly at random from valid moves.
4.3 Software
We implemented AlphaViT, AlphaViD, the opponents, and the board games in Python, using NumPy for linear algebra and PyTorch for deep learning components. The source code is available on GitHub at https://github.com/KazuhisaFujita/AlphaViT for reproducibility and extension of this work.
5 Results
In this results, we investigate the performance and characteristics of the proposed AI game-playing agents: AlphaViT, AlphaViD, and AlphaVDA. These agents have been trained on different games (Connect4, Gomoku, and Othello) with varying board sizes (small and large). These agents employ diverse architectures in terms of the number of encoder layers to examine how the numbers of encoder layers choices impact the learning outcomes and game skill of the AI agents.
The architectures of these agents vary primarily by the number of encoder layers, which influences their capacity to learn from the game environments. The table LABEL:tab:num_params shows the number of parameters for each agent configuration. AlphaViT, AlphaViD, and AlphaVDA are tested with different numbers of encoder layers, denoted by ‘L’ followed by a number (e.g., L1, L4, L5, L8). The number of parameters ranges from 11.2 million to 19.9 million, increasing with the number of encoder layers. For comparison, the AlphaZero agent, which serves as a baseline, has 7.1 million parameters.
Table 1: Encoder Layer Variations and Parameter Sizes in AI Agents
AI agent num of encoder layers num of parameters
AlphaViT L4 4 11.2M
AlphaViD L1 1 11.5M
AlphaVDA L1 1 11.3M
AlphaViT L8 8 19.6M
AlphaViD L5 5 19.9M
AlphaVDA L5 5 19.8M
AlphaZero - 7.1M
Each AI agent is trained on specific games with varying board sizes. The table 2 categorizes the agents based on the games which they are trained on and the board sizes used during training. The first group consists of agents trained on a single game with small board size, denoted as SB. The second group includes agents trained on a single game with large board size, denoted as LB. Finally, the third group comprises agents trained on multiple games, including Connect4, Gomoku, and Othello, with large board sizes, denoted as Multi. The agents in the third group are trained simultaneously the three games. They can play these three game using only one DNN. In other words, they are not specialized for one specific game. This diversity in training regimes allows us to evaluate the agents’ adaptability and generalization capabilities across different game domains.
Table 2: Board Size and Game Variations in AI Agent Training
AI agents Game Board size
AlphaViT SB, AlphaViD Small SB, AlphaVDA Small SB one specific game Small
AlphaViT LB, AlphaViD Small LB, AlphaVDA Small LB one specific game Large
AlphaViT Multi, AlphaViD Multi, AlphaVDA Multi Connect4, Gomoku, Othello Large
5.1 Elo rating of board game playing algorithms
Table 3 shows the Elo ratings of various AI agents for different games and board sizes. Elo ratings serve as a standard measure of relative skills in two-player games and allow a systematic comparison of the performance of each AI agent. The agents include variations of the proposed methods (AlphaViT, AlphaViD, and AlphaVDM) trained on large boards (LB), small boards (SB), and multiple board sizes (Multi), as well as other AI agents, including AlphaZero, Monte Carlo Tree Search (MCTS) with different numbers of simulations, Minimax, and a Random agent. The proposed methods and AlphaZero are trained through 1000 iterations. The Elo rating is calculated through 50 round-robin tournaments between the agents. The Elo ratings of all agents are initialized to 1500. The details of the Elo rating calculation are given in Appendix 5.
AlphaViT L4 LB shows strong performance, achieving Elo ratings of 1824 in Connect4, 1835 in Gomoku, and 2125 in Othello. However, for Connect4 and Gomoku AlphaViT is outperformed by AlphaZero. Interestingly, despite not being specifically trained on small boards, AlphaViT L4 LB performs as well or better than Minimax and MCTS in Connect4 5x4, Gomoku 6x6, and Othello 6x6. This suggests that AlphaViT L4 SB efficiently utilizes the knowledge obtained from large board training.
AlphaViT L4 SB shows Elo ratings of 1732 in Connect4 5x4, 1764 in Gomoku 6x6 and 1969 in Othello 6x6. In these games AlphaViT L4 SB performed as well as or better than AlphaZero. For Gomoku and Othello, it outperformed Minimax and MCTS even though it is not trained on large board games. This result suggests that AlphaViT L4 SB effectively transfers knowledge from small board training to large boards.
AlphaViT L4 Multi, trained simultaneously on Connect4, Gomoku and Othello, showed strong results in these games, achieving Elo ratings of 1700 in Connect4, 1969 in Gomoku and 1939 in Othello. Although it is weaker than AlphaZero in Connect4 and Gomoku, AlphaViT L4 Multi outperformed AlphaViT L4 LB (trained on Gomoku only) in Gomoku. In addition, AlphaViT L4 Multi shows comparable performance to both AlphaViT L4 SB and AlphaZero in Gomoku 6x6.
AlphaViD L1 LB performs slightly worse than AlphaViT in the games with larger boards. In Gomoku 6x6 and Othello 6x6, AlphaViD L1 LB outperformed both Minimax and MCTS100, suggesting that it effectively transferred knowledge gained from training on larger boards. AlphaViD L1 SB achieved Elo ratings of 1742 in Connect4 5x4, 1785 in Gomoku 6x6 and 1852 in Othello 6x6, with performance comparable to AlphaZero and AlphaViT. AlphaViD L1 Multi showed comparable performance to AlphaViT in Connect4, but was weaker in Gomoku and Othello.
Similarly, AlphaVDA L1 LB is weaker than AlphaViT for the games with large boards. AlphaVDA L1 LB outperformed Minimax and MCTS100 for all games with small board sizes. AlphaVDA L1 SB shows Elo ratings of 1770 in Connect4 5x4, 1871 in Gomoku 6x6 and 1833 in Othello 6x6, with performance comparable to AlphaZero, AlphaViT and AlphaViD. AlphaVDA L1 Multi performed better than AlphaViT in Connect4, but was outperformed by AlphaViT in both Gomoku and Othello.
The Elo ratings presented in this table serve as a baseline for the subsequent experiments described in the following sections.
Table 3: Elo Ratings of AI Agents Across Different Games and Board Sizes
Game Connect4 Connect4 5x4 Gomoku Gomoku 6x6 Othello Othello 6x6
AlphaViT L4 LB 1824 1477 1835 1572 2125 1820
AlphaViD L1 LB 1755 1472 1724 1532 1856 1581
AlphaVDA L1 LB 1757 1506 1548 1244 1745 1336
AlphaViT L4 SB 1169 1732 1564 1764 1487 1969
AlphaViD L1 SB 1187 1742 1503 1785 1128 1852
AlphaVDA L1 SB 1213 1770 973.5 1871 1131 1833
AlphaViT L4 Multi 1700 1374 1969 1785 1939 1136
AlphaViD L1 Multi 1734 1295 1538 1654 1695 1378
AlphaVDA L1 Multi 1807 1321 1524 1194 1564 953.2
AlphaZero 2003 1767 2279 1769 1889 1769
Minimax 1033 1337 1484 1403 1127 1403
MCTS100 1260 1512 1153 1414 1192 1414
MCTS400 1550 1581 1169 1675 1365 1675
Random 725.8 1057 697.0 836.2 741.5 838.6
6 Conclusion and discussion
We propose AlphaViT, AlphaViD, and AlphaVDA, which are game-playing AI agents designed to address the limitations of AlphaZero using ViT. Unlike AlphaZero, which is limited to fixed board sizes, these proposed methods can effectively handle variations in board size, demonstrating flexibility and adaptability in their gameplay across different board sizes and game types. In addition, we showed that AlphaViT, AlphaViD, and AlphaVDA can simultaneously train and play multiple games, such as Connect4, Gomoku, and Othello, using a single model trained on all games. This ability to generalize across games with a single model represents a significant advance over traditional game-specific AI models.
The results of our experiments show that AlphaViT and AlphaViD outperform baseline methods such as Minimax and MCTS in most configurations. Although AlphaZero still achieves the highest Elo ratings in some cases, especially in games with larger boards, the proposed agents demonstrate competitive performance, especially in Othello, where AlphaViT approaches and even exceeds AlphaZero’s Elo rating in specific configurations. Furthermore, the multi-game versions of AlphaViT and AlphaViD perform on par with their single-game counterparts, further highlighting the ability of these architectures to generalize and adapt across different board sizes.
AlphaViT and AlphaViD showed strong adaptability across different games and board sizes. In particular, agents trained on single games often demonstrate performance comparable to AlphaZero, even when playing on board sizes for which they were not explicitly trained. This suggests effective knowledge transfer between different board sizes, mirroring human learning processes in which skills from simpler game variants (e.g., 9x9 Go) can be applied to more complex versions (19x19 Go). This similarity to human learning patterns suggests that such training paradigms may be beneficial for AI development, particularly in the context of multitask learning.
A comparison of the three proposed architectures shows that AlphaViT performs slightly better than AlphaViD and AlphaVDA despite having the same number of parameters. This may be due to the fewer transformer encoder layers in AlphaViD and AlphaVDA, which rely on a more complex architecture with both encoder and decoder layers. AlphaViT’s simpler architecture, consisting only of encoder layers, may benefit from having more layers dedicated to learning, resulting in more efficient performance in certain games. However, this simplicity limits AlphaViT’s flexibility because its output size is fixed to the number of input tokens, reducing its applicability to games beyond the classic board games tested here. In contrast, AlphaViD’s inclusion of a decoder layer allows it to adjust the size of its policy vector dynamically, providing greater adaptability to games with different action spaces. This architectural flexibility will make AlphaViD more versatile for handling complex games or environments with continuous action spaces.
Future work will explore the application of these architectures to a broader range of games, including those with more complex rules and non-deterministic elements. In addition, we will extend the flexibility of ViT to other deep reinforcement learning methods, such as deep Q-network, and develop a game AI agent that can play more flexibly, including computer games.
Appendix
1 AlphaZero
AlphaZero consists of a deep neural network (DNN) and Monte Carlo tree search (MCTS), as shown in Fig. 1. The DNN receives an input representing the current state of the board and the current player. It then outputs the estimated state value and the move probability. MCTS determines the best move based on the value and move probability. AlphaViT, AlphaViD, and AlphaVDA adopt this same fundamental structure, employing an identical decision-making process to select their next moves.
1.1 Deep neural network in AlphaZero
AlphaZero’s deep neural network (DNN) predicts the value $v(s)$ and the move probability $\bm{p}$ with components $p(a|s)$ for each action $a$, given a state $s$. In a board game context, $s$ and $a$ represent the board state and the move, respectively. The DNN receives an input representing the current board state and the current player’s disc color. Fig. LABEL:fig:network illustrates the DNN architecture, which consists of a Body (residual blocks) and a Head (value and policy heads). The value head outputs the estimated state value $v(s)$, while the policy head produces the move probabilities $\bm{p}$.
The input to the DNN is an $M\times N\times(2T+1)$ image stack that contains $12T+$ 1 binary feature planes of size $M\times N$. Here, $M\times N$ refers to the board size, and $T$ is the number of histories. The first $T$ feature planes represent the occupancy of the player’s discs, with a feature value of 1 indicating that a disc occupies the corresponding cell, and 0 indicating otherwise. Similarly, the following $T$ feature planes represent the occupancy of the other players’ discs. The last feature plane represents the disc color of the current player, with the disc colors of the first and second players being represented by 1 and -1, respectively.
1.2 Monte Carlo tree search in AlphaZero
This subsection provides an explanation of the Monte Carlo Tree Search (MCTS) algorithm used in AlphaZero. Each node in the game tree represents a game state, and each edge $(s,a)$ represents a valid action from that state. The edges store a set of statistics: $\{N(s,a),W(s,a),Q(s,a),P(s,a)\}$, where $N(s,a)$ is the visit count, $W(s,a)$ is the cumulative value, $Q(s,a)=W(s,a)/N(s,a)$ is the mean value, and $P(s,a)$ is the move probability.
The MCTS for AlphaZero consists of four steps: Select, Expand and Evaluate, Backup, and Play. A simulation is defined as a sequence of Select, Expand and Evaluate, and Backup steps, repeated $N_{\mathrm{sim}}$ times. Play is executed after $N_{\mathrm{sim}}$ simulations.
In Select, the tree is searched from the root node $s_{\mathrm{root}}$ to the leaf node $s_{L}$ at time step $L$ using a variant of the PUCT algorithm. At each time step $t<L$, the selected action $a_{t}$ has a maximum score, as described by the following equation:
$a_{t}=\mathop{\rm arg~{}max}\limits_{a}(Q(s_{t},a)+C_{\mathrm{puct}}P(s_{t},a)\frac{\sqrt{N(s_{t})}}{1+N(s_{t},a)}),$ (8)
where $N(s_{t})$ is the number of parent visits and $C_{\mathrm{puct}}$ is the exploration rate. In this study, $C_{\mathrm{puct}}$ is constant, whereas in the original AlphaZero, $C_{\mathrm{puct}}$ increases slowly with search time.
In Expand and Evaluate, the DNN evaluates the leaf node and outputs $v(s_{l})$ and $\bm{p}_{a}(s_{l})$. If the leaf node is a terminal node, $v(s_{L})$ is the color of the winning player’s disc. The leaf node is expanded and each edge $(s_{L},a)$ is initialized to $\{N(s_{L},a)=0,W(s_{L},a)=0,Q(s_{L},a)=0,P(s_{L},a)=p_{a}\}$.
In Backup, the visit counts and values are updated in a backward pass through each step, $t\leq L$. The visit count is incremented by 1, $N(s_{t},a_{t})\leftarrow N(s_{t},a_{t})+1$, and the cumulative and average values are updated, $W(s_{t},a_{t})\leftarrow W(s_{t},a_{t})+v$, $Q(s_{t},a)\leftarrow W(s_{t},a_{t})/N(s_{t},a_{t})$.
Finally, in Play, AlphaZero selects the action corresponding to the most visited edge from the root node.
2 Training
AlphaViT, AlphaViD, AlphaVDA, and AlphaZero use a common training scheme. The training process consists of three main components: Self-play, Augmentation, and Learning, which are iterated $N_{\mathrm{iter}}$ times. This training algorithm is a modified version of the original AlphaZero, adapted to allow training on a single computer.
During the Self-play phase, the AI agent plays against itself $N_{\mathrm{self}}$ times. For the first $T_{\mathrm{opening}}$ turns, actions are stochastically selected from valid moves based on the softmax policy:
$p(a\mid s)=\exp(N(s,a)/\tau)/\Sigma_{b}\exp(N(s,b)/\tau),$ (9)
where $\tau$ is a temperature parameter that controls the exploration. This stochastic exploration enables the agent to explore new and potentially better actions. After $T_{\mathrm{opening}}$ the most visited action is selected. Through Self-play, we collect board states, winners, and search probabilities. The search probabilities represent the probabilities of selecting valid moves at the root node in MCTS.
In the Augmentation phase, the dataset derived from Self-play is augmented by introducing symmetries specific to the game variant (e.g., two symmetries for Connect4 and eight for Othello and Gomoku). This augmented data are added to a queue with a capacity of $N_{\mathrm{queue}}$ states to form the training dataset.
For the first learning iteration, the training data queue is filled with data generated by self-play using MCTS100 and augmented them. In subsequent iterations, new data generated by Self-play are added to the training data queue. To learn multiple games simultaneously, we prepare a separate training data queue for each game.
During the Learning phase, the deep neural network (DNN) is trained using mini-batch stochastic gradient descent with $N_{\mathrm{batch}}$ batch size and $N_{\mathrm{epochs}}$ epochs. The optimization process includes momentum and weight decay. The loss function $l$ combines the mean squared error between the predicted value $v$ and the winner’s disc color $c_{\mathrm{win}}$, and the cross-entropy loss between the search probabilities $\bm{\pi}$ and the predicted move probabilities $\bm{p}$:
$l=(c_{\mathrm{win}}-v)^{2}-\bm{\pi}^{\mathrm{T}}\log\bm{p}.$ (10)
To train multiple games simultaneously, mini-batches are generated from the respective training data queue of each game. During the Learning phase, mini-batches are sampled from these individual queues and used to update the DNN. For example, when an agent simultaneously trains Connect4, Gomoku, and Othello, the mini-batch of Connect4, the mini-batch of Gomoku, and the mini-batch of Othello are used orderly for update.
3 Parameters
The hyperparameters for AlphaViT, AlphaViD, AlphaVDA, and AlphaZero are detailed in Table LABEL:tab. The weight decay and momentum values are consistent with those specified in the AlphaZero pseudocode (Silver et al., 2018). All other parameters for AlphaZero are consistent with the previous implementation Fujita (2022). The hyperparameters for the other models were carefully hand-tuned to optimize their performance.
Table LABEL:tab:param_alphavit_game lists the game-specific hyperparameters for AlphaViT, ALphaViD, AlphaVDA, and AlphaZero. The number of MCTS simulations ($N_{\mathrm{sim}}$) ranges from 200 to 400, depending on the game and board size. The number of self-play games per iteration is set to 30 for Connect4 variants and 10 for Gomoku and Othello variants. The opening phase ($T_{\mathrm{opening}}$) specifies the number of initial moves using softmax decision-making with a temperature parameter ($\tau$) that is adjusted based on the game and board size. These hyperparameters were also carefully hand-tuned.
Table A1: Hyperparameters of AlphaViT, AlphaViD, AlphaVDA, and AlphaZero
parameter AlphaViT AlphaViD AlphaVDA AlphaZero
Num iterations 1000 1000 1000 1000
$C_{\mathrm{puct}}$ 1.25 1.25 1.25 1.25
$\epsilon$ 0.2 0.2 0.2 0.2
$T$ 1 1 1 1
$N_{\mathrm{queue}}$ 100000 100000 100000 100000
$N_{\mathrm{epoch}}$ 1 1 1 1
batch size 1024 1024 1024 1024
Learning rate 0.01 0.01 0.01 0.01
Momentum 0.9 0.9 0.9 0.9
Weight decay 0.0001 0.0001 0.0001 0.0001
patch size 5 5 5 -
stride of a patch 1 1 1 -
Num of encoder layers $L$ $L$ $L$
Embedding size of encoder 512 512 512 -
forward size of encoder 1024 1024 1024 -
Num of encoder head 16 16 16 -
Num of decoder layers - 1 1 -
Embedding size of decoder - 512 512 -
forward size of encoder - 1024 1024 -
Num of decoder head - 16 16 -
Action token size - - 256 -
Num of residual blocks - - - 3
Kernel size - - - 3
Number of filters - - - 256
Table A2: Game-specific Hyperparameters for AlphaViT, AlphaViD, AlphaVDA, and AlphaZero
Connect4 Connect4 5x4 Gomoku Gomoku 66 Othello Othello 66
Num of simulations 200 200 400 200 400 200
Num of self-play 30 30 10 10 10 10
$T_{\mathrm{opening}}$ 6 4 8 6 6 4
$\tau$ 100 100 40 20 80 40
4 Minimax algorithm
The minimax algorithm is a fundamental game-tree search technique that determines the optimal action by evaluating the best possible outcome for the current player. Each node in the tree contains a state, player, action, and value. The algorithm creates a game tree with a depth of $N_{\mathrm{depth}}=3$. The root node corresponds to the current state and minimax player. Next, the states corresponding to leaf nodes are evaluated. Then, the algorithm propagates the values from the leaf nodes to the root node. If the player corresponding to the node is the opponent, the value of the node is the minimum value of its child nodes. Otherwise, the value of the node is the maximum value of its child nodes. Finally, the algorithm selects the action corresponding to the root’s child node with the maximum value. The evaluation of leaf nodes is tailored to each game. For the Connect4 variants, the values of the connections of two and three same-colored discs are $R\times c_{\mathrm{disc}}c_{\mathrm{minimax}}$ and $R^{2}\times c_{\mathrm{disc}}c_{\mathrm{minimax}}$, respectively, where $R$ is the base reward, and $c_{\mathrm{disc}}$ and $c_{\mathrm{minimax}}$ are the colors of the connecting discs and the minimax player’s disc, respectively. The value of a node is the sum of the values of all connections on the corresponding board. The terminal nodes have a value of $R^{3}c_{\mathrm{win}}c_{\mathrm{minimax}}$, where $c_{\mathrm{win}}$ is the color of the winner, and $R=100$.
For the Gomoku variants, the values of the connections of two, three, and four same-colored discs are $R\times c_{\mathrm{disc}}c_{\mathrm{minimax}}$, $R^{2}\times c_{\mathrm{disc}}c_{\mathrm{minimax}}$, and $R^{3}\times c_{\mathrm{disc}}c_{\mathrm{minimax}}$, respectively. The value of a node is the sum of the values of all connections on the corresponding board. The terminal nodes have values of $R^{4}c_{\mathrm{win}}c_{\mathrm{minimax}}$ and $R=100$.
For the Othello variants, the value of a node is calculated using the following equation:
$E=\sum_{x}\sum_{y}v(x,y)o(x,y)c_{\mathrm{minimax}},$ (11)
where $v_{(}x,y)$ is the value of cell $(x,y)$ and $o(x,y)$ is the occupancy of cell $(x,y)$. For 6x6 Othello and 8x8 Othello, the minimax algorithm evaluates each cell using Eq. 12 and 13, respectively. $o(x,y)$ is $1$, $-1$, and 00 if cell $(x,y)$ is occupied by the first player’s disc, the second player’s disc, and empty, respectively. For the Othello variants, the algorithm expands the tree to the terminal nodes after the last six turns. The terminal nodes have a value of $E_{\mathrm{end}}=1000c_{\mathrm{win}}c_{\mathrm{minimax}}$.
$v_{6\times 6}=\left(\begin{matrix}30&-5&2&2&-5&30\\ -5&-15&3&3&-15&-5\\ 2&3&0&0&3&2\\ 2&3&0&0&3&2\\ -5&-15&3&3&-15&-5\\ 30&-5&2&2&-5&30\\ \end{matrix}\right).$ (12)
$v_{8\times 8}=\left(\begin{matrix}120&-20&20&5&5&20&-20&120\\ -20&-40&-5&-5&-5&-5&-40&-20\\ 20&-5&15&3&3&15&-5&20\\ 5&-5&3&3&3&3&-5&5\\ 5&-5&3&3&3&3&-5&5\\ 20&-5&15&3&3&15&-5&20\\ -20&-40&-5&-5&-5&-5&-40&-20\\ 120&-20&20&5&5&20&-20&120\\ \end{matrix}\right).$ (13)
5 Elo rating
Elo rating is a widely used metric for evaluating the relative skill levels of players in two-player games. It allows us to estimate the probability of one player defeating another based on their current ratings. Given two players A and B with Elo ratings $e(A)$ and $e(B)$, respectively, the probability that player A will defeat player B, denoted $p(A\text{ defeats }B)$, is calculated using the following formula:
$p(A~{}\rm{defeats}~{}B)=1/(1+10^{(e(B)-e(A))/400}).$ (14)
After a series of $N_{G}$ games between players A and B, player A’s Elo rating is updated to a new value $e^{\prime}(A)$ based on their performance:
$e^{\prime}(A)=e(A)+K(N_{\mathrm{win}}-N_{G}\times p(A~{}\rm{defeats}~{}B)),$ (15)
where $N_{\mathrm{win}}$ is the number of times player A has won, and $K$ is a factor that determines the maximum rating adjustment after a single game. In this study, $K=8$.