### v2

# AlphaViT: A Flexible Game-Playing AI for Multiple Games and Variable Board Sizes

Abstract
This paper presents novel game-playing AI agents based on the AlphaZero framework, enhanced with Vision Transformer (ViT): AlphaViT, AlphaViD, and AlphaVDA. These agents are designed to play multiple board games of various sizes using a single network with shared weights, thereby overcoming AlphaZero’s limitation of fixed-board-size constraints. AlphaViT employs only a transformer encoder, whereas AlphaViD and AlphaVDA incorporate both transformer encoders and decoders. In AlphaViD, the decoder processes outputs from the encoder, whereas AlphaVDA uses a learnable embeddings as the decoder input. The additional decoder layers in AlphaViD and AlphaVDA provide flexibility to adapt to various action spaces and board sizes. Experimental results show that the proposed agents, trained on either individual games or multiple games simultaneously, consistently outperform traditional algorithms such as Minimax and Monte Carlo Tree Search and approach the performance of AlphaZero, despite using a single deep neural network (DNN) with shared weights. In particular, AlphaViT shows strong performance across all tested games. Furthermore, fine-tuning the DNN using pre-trained weights from small-board games accelerates convergence and improves performance, particularly in Gomoku. Interestingly, simultaneous training on multiple games yields performance comparable to, or even surpassing, single-game training. These results indicate the potential of transformer-based architectures to develop more flexible and robust game-playing AI agents that excel in multiple games and dynamic environments.
1 Introduction
Artificial intelligence (AI) has advanced remarkably in recent years, demonstrating its potential across a wide range of applications. One area where AI has excelled is in mastering board games, often outperforming top human players. Notable achievements include AI systems that defeat humans in games, such as checkers, chess (Campbell et al., 2002), and Othello (Buro, 1997). A significant turning point occurred in 2016 when AlphaGo (Silver et al., 2016), an AI designed specifically for the Go game, defeated the world’s top players. After this, AlphaZero (Silver et al., 2018) was introduced, demonstrating its capability to master various board games, including Chess, Shogi, and Go. This further highlights AI’s superhuman skills of AI in this domain.
Despite these successes, many current game-playing AI agents suffer from a fundamental limitation: they are typically designed for only a single specific game and cannot play other games. Even when the game rules remain consistent, these agents cannot handle variations in board size. In contrast, humans can easily switch between different board sizes. For example, beginners in Go often start practicing on smaller boards (e.g., $9\times 9$) before progressing to larger boards (e.g., $19\times 19$). However, AI agents like AlphaZero, which are designed for a single specific game and fixed board size, require substantial reprogramming to accommodate such changes.
For AlphaZero, this limitation arises from its deep neural network (DNN) architecture, which requires a fixed input size. AlphaZero’s DNN consists of residual blocks and multilayer perceptrons (MLPs). These components are designed for a specific input size. The output size of the residual blocks varies with changes in input size, creating inconsistencies with the expected MLP input size. Consequently, AlphaZero cannot function properly even with small variations in board size.
To address this limitation, this study proposes replacing residual blocks in the AlphaZero framework with Vision Transformer (ViT) (Dosovitskiy et al., 2021). ViT is an image-classification DNN based on the transformer architecture. ViT divides an image into patches and infers a class of the image from the patches. A key advantage of ViT is its flexibility in handling images of various sizes. This flexibility enables the AlphaZero framework to adapt to various games and board sizes.
This paper presents game-playing agents capable of handling multiple games and variable board sizes using a single DNN. These agents, named AlphaViT, AlphaViD (AlphaViT with a transformer decoder layer), and AlphaVDA (AlphaViD with learnable action embeddings), are based on the AlphaZero framework. The agents predict the value of a game state and policy using a DNN, while their decisions are made using Monte Carlo Tree Search (MCTS). Our computational experiments show that the proposed agents can be simultaneously trained on three games (Connect4, Gomoku, and Othello) using a single DNN with shared weights. Moreover, the proposed agents outperform traditional algorithms such as Minimax and MCTS across various games while approaching the performance of AlphaZero, whether trained on a single game or simultaneously trained on multiple games.
Portions of this text were previously published as part of a preprint (https://arxiv.org/abs/2408.13871).
2 Related work
Game-playing AI agents have reached superhuman performance levels in traditional board games such as Checkers (Schaeffer et al., 1993), Othello (Buro, 1997, 2003), and Chess (Campbell, 1999; Hsu, 1999; Campbell et al., 2002). In 2016, AlphaGo (Silver et al., 2016), a Go-playing AI, has defeated the world’s top Go players, marking the first superhuman-level performance in Go. AlphaGo relies on supervised learning from a large database of expert human moves and self-play data. Subsequently, AlphaGo Zero (Silver et al., 2017) has defeated AlphaGo, relying only on self-play data. In 2018, Silver et al. (2018) have proposed AlphaZero, which has no restrictions on playable games. AlphaZero has outperformed other superhuman-level AIs in Go, Shogi, and Chess. Interestingly, AlphaZero’s capabilities extend beyond traditional two-player perfect information games, with research exploring its potential in more complex scenarios. For example, Hsueh et al. (2018) have shown AlphaZero’s potential in nondeterministic games. Other extensions include handling continuous action spaces (Moerland et al., 2018) and support for multiplayer games (Petosa and Balch, 2019). However, AlphaZero cannot play multiple games simultaneously and handle games with the same rules but different board sizes using a single DNN with shared weights. This limitation is due to the use of the DNN in AlphaZero for policy and value estimations. AlphaZero’s DNN consists of residual blocks and MLPs. While the residual blocks can process input images of various sizes, the MLPs fix their input sizes. Consequently, the DNN is not sufficiently flexible to accommodate variations in board size.
Researchers have improved the AlphaZero framework to overcome its limitations. Wu (2019) has tackled some of these limitations by improving the efficiency of AlphaZero-like training in Go. His model introduces techniques such as playout cap randomization and policy target pruning, which significantly accelerate self-play learning. Wu has used global pooling layers to standardize varying board sizes to a fixed size, enabling the model to estimate values using MLPs effectively. Furthermore, Wu has exclusively utilized convolutional layers for policy estimation, thereby avoiding the need for MLPs for policy estimation. This design enables seamless handling of varying board sizes while efficiently calculating the policy. This innovation marks a significant step towards creating more flexible AI agents capable of adapting to different board configurations. Similarly, Soemers et al. (2023) have explored transfer learning across various board games, employing fully convolutional networks with global pooling to enable effective transfer between games with different board sizes, shapes, and rules. Their approach has demonstrated the potential of convolutional architectures, enhanced with global pooling, to generalize across various game scenarios.
To address the limitation of the AlphaZero framework, we focus on the transformer architecture and solve the problem by integrating transformer architecture into the AlphaZero framework. The transformer architecture, initially developed for natural language processing (Vaswani et al., 2017), has shown remarkable effectiveness in various domains. The transformer architecture has also been successfully applied to image-processing tasks. Transformer-based models have achieved exceptional performance in various image-related tasks, including image classification (Dosovitskiy et al., 2021), semantic segmentation (Xie et al., 2021), video classification (Li et al., 2022), and video captioning (Zhao et al., 2022). ViT, introduced by Dosovitskiy et al. (Dosovitskiy et al., 2021), is a remarkable example of a transformer-based model for image processing. ViT has achieved state-of-the-art performance in image classification at the time of its introduction. A key feature of ViT is its independence from the input image size (Dosovitskiy et al., 2021). Unlike convolutional neural networks, which require fixed-size inputs, ViT can process images of various sizes by dividing them into fixed-size patches and treating each patch as a token in the transformer architecture. This flexibility makes ViT highly adaptable and efficient when handling different image sizes. By incorporating the ViT architecture into the AlphaZero framework, we extend its capability to handle various games with different board sizes, enhancing its flexibility.
3 AlphaViT, AlphaViD, and AlphaVDA
AlphaZero’s game-playing capability is limited to games with specific board sizes and rules used during training. This limitation stems from AlphaZero’s deep neural network (DNN), which consists of residual blocks followed by multilayer perceptrons (MLPs) for value and policy computations. The MLPs receive input from the final residual block and assume a fixed input size. Consequently, AlphaZero’s performance is restricted to games with predetermined board size.
To address this limitation, AlphaViT, AlphaViD, and AlphaVDA are developed as game-playing AI agents based on the AlphaZero framework but using Vision Transformer (ViT) architecture. These agents use a combination of a DNN and MCTS (Fig. 1). The DNN receives the board state, and outputs the value estimates and the move probabilities (policies). The MCTS searches a game tree using the estimated value and move probability. By incorporating ViT instead of residual blocks, AlphaViT, AlphaViD, and AlphaVDA can overcome the limitations of AlphaZero and play games with varying board sizes and rules.
An overview of AlphaViT’s DNN architecture is shown in Fig. 2. The DNN in AlphaViT is based on ViT, which has no input-size limitation and can classify images even if the input image size differs from the training image size. This flexibility enables AlphaViT to play games with different board sizes using the same network. In AlphaViT, the game boards are fed into ViT. Initially, these inputs are transformed into patch embeddings through a convolutional layer. Using the convolution layer allows for easy adjustment of patch division parameters such as patch size, stride, and padding. The final outputs of the encoder layer are used to compute the value and the move probabilities.
AlphaViD and AlphaVDA incorporate both transformer encoder and decoder layers for calculating value and move probabilities, as shown in 3. In AlphaViD, the decoder layer receives the input derived from the output of the encoder layer. In contrast, AlphaVDA employs learnable action embeddings as inputs to the decoder layer. The final outputs of the encoder layer are used to compute the value, while the final outputs of the decoder layer are used to compute the move probabilities.
Importantly, AlphaViT, AlphaViD, and AlphaVDA can play any game that AlphaZero can because they employ the same game-playing algorithm as AlphaZero.
Figure 1: AlphaViT, AlphaViD, and AlphaVDA follow the same procedural framework as AlphaZero. They take the board state and a current player as inputs and determine the next move using Monte Carlo Tree Search (MCTS). MCTS explores the game tree using the value and policy probabilities provided by the DNN.
Figure 2: Overview of the AlphaViT model architecture. An input board state is divided into fixed-size patches and linearly embedded by a convolutional layer. Position embeddings are added to the patch embeddings. The resulting sequence of the patch embeddings is appended with a value embedding, a game embedding, and a pass embedding. The sequence is fed into a transformer encoder. Finally, a multilayer perceptron (MLP) estimates the value and move probabilities by taking the outputs from the last layer of the transformer encoder as its input.
Figure 3: Architecture of AlphaViD and AlphaVDA models. An input board state is divided into fixed-size patches and linearly embedded using a convolutional layer. Position embeddings are added to the resulting patch embeddings to preserve spatial information. The sequence of patch embeddings is appended with a value embedding and a static game embedding. This input sequence is fed into a transformer encoder. The estimated value is obtained from a multilayer perceptron (MLP) that takes the output corresponding to the value embedding from the last layer of the transformer encoder. The output embeddings are input to the transformer decoder. For AlphaViD, the input embeddings for the transformer decoder are generated from the output embeddings corresponding to the board state by a fully connected layer. For AlphaVDM, the input embeddings of the transformer decoder are generated from learnable action embeddings. Finally, the move probabilities are produced by another MLP, which takes the output of the last layer of the transformer decoder.
3.1 Architectures of DNNs
AlphaViT
Given a game state $s$, AlphaViT’s DNN predicts the value $v(s)$ and the move probability distribution $\bm{p}(s)$, which includes $p(a|s)$ for each action $a$. In a board game, $s$ represents the board state, and $a$ denotes a move.
The input to the DNN is an $H\times W\times(2T+1)$ image stack ${\mathbf{x}}\in\mathbb{R}^{H\times W\times(2T+1)}$, consisting of $12T+$ 1 binary feature planes of size $H\times W$. Here, $H$ and $W$ are the dimensions of the board, and $T$ is the number of history planes. The first $T$ feature planes represent the occupancies of the first player’s discs, where a feature value of 1 means that a disc occupies the corresponding cell and, and 0 means it does not. The following $T$ feature planes represent the occupancies of the second player’s discs. The final feature plane represents the color of the current player’s disc, where 1 denotes the first player and -1 denotes the second player.
The convolutional layer divides the image stack ${\mathbf{x}}$ into $P\times P$ patches with stride $s$ and padding $p$, and generates the patch embeddings ${\mathbf{z}}_{0}$. The patch size $P$ corresponds to the kernel size of the convolutional layer. The sequence of patch embeddings ${\mathbf{z}}_{0}$ is expressed as
${\mathbf{z}}_{0}=[{\mathbf{x}}^{1}_{p}{\mathbf{E}};...;{\mathbf{x}}^{i}_{p}{\mathbf{E}};...;{\mathbf{x}}^{N_{p}}_{p}{\mathbf{E}}],$ (1)
where ${\mathbf{x}}^{i}_{p}\in\mathbb{R}^{P^{2}(2T+1)}$ is the $i$th flattened 2D patch, and ${\mathbf{E}}$ is a trainable embedding tensor with the shape $(P^{2}(2T+1),D)$. The kernel of the convolutional layer acts as the tensor ${\mathbf{E}}$, which maps each patch to a $D$-dimensional embedding space. Here, $N_{p}$ is the number of embeddings, and $D$ is the embedding size.
To retain position information, learnable 2D position embeddings ${\mathbf{E}}_{\mathrm{pos}}$ are added to the patch embeddings. These positional embeddings are scaled according to the board size and hyperparameters to match the patch embeddings ${\mathbf{x}}_{0}$. The position embeddings are incorporated into the token embeddings ${\mathbf{z}}_{0}$ as follows:
${\mathbf{z}}_{0}={\mathbf{z}}_{0}+E_{\mathrm{pos}}.$ (2)
The output size of the transformer encoder layer is determined by the number of input embeddings. For Gomoku, where the action space is $HW$, AlphaViT requires $N_{p}=HW$ embeddings. To achieve this, we set the patch size $P=2k+1$, where $k$ is a positive integer, stride $s=1$, and padding $\mathrm{pad}=\lfloor P/2\rfloor$. In contrast, for Othello, the action space is $HW+1$ to include the pass move. Using the same parameters as Gomoku ($P=2k+1$, $s=1$, and $\mathrm{pad}=\lfloor P/2\rfloor$), AlphaViT requires one additional embedding for the pass move. To address this, we introduce a learnable pass token $\bm{x}_{\mathrm{pass}}$ and append it to the input embeddings ${\mathbf{z}}_{0}$. To estimate the board value, we prepend a learnable value embedding $\bm{x}_{\mathrm{value}}$. Additionally, to enable AlphaViT to recognize different game types, we incorporate a static game embedding $\bm{x}_{\mathrm{game}}$, represented using one-hot encoding. These embeddings are appended to the input embeddings ${\mathbf{z}}_{0}$, resulting in
${\mathbf{z}}_{0}=[\bm{x}_{\mathrm{value}};\bm{x}_{\mathrm{game}};\bm{x}^{1}_{p}\bm{\mathrm{E}};...;\bm{x}^{N_{p}}_{p}\bm{\mathrm{E}};\bm{x}_{\mathrm{pass}}].$ (3)
Position embeddings are added before appending the pass, value, and game embeddings. This approach ensures that the position embeddings can be scaled independently without being affected by embeddings that do not inherently contain positional information.
The sequence $\bm{z}_{0}$ is fed into the transformer encoder, which consists of $L$ transformer encoder layers. The output of the final encoder layer ${\mathbf{z}}_{L}$ consists of $M\times N+3$ vectors. The first vector $z^{0}_{L}$ derived from the value embedding is processed by the value head implemented as a multilayer perceptron (MLP) denoted as $\text{MLP}_{v}$. This head estimates the value $v$ using the following equation:
$v=\tanh(\mathrm{NLP}_{v}(\mathrm{LN}(z^{0}_{L}))),$ (4)
where LN represents layer normalization. The tanh activation constrains the value within the range $(-1,1)$, representing the winning probability.
The vectors $z^{2}_{L},...,x^{WH+2}_{L}$, derived from the board patches and the pass embedding, are processed by the policy head, which is implemented as another MLP ($\mathrm{NLP}_{p}$). The policy head outputs the move probability $p(m,n)$ for each board position $(m,n),0\leq m\leq W,0\leq n<H$ using the sigmoid function. The policy head outputs a vector $\mathbb{R}^{H\times W+1}$ representing the move probabilities
$p(m,n)=\mathrm{Sigmoid}(\mathrm{MLP}_{p}(\mathrm{LN}(z^{m*H+n+2}_{L})))(m,n).$ (5)
For Othello, $p(W,0)$ corresponds to the probability of a pass action. In contrast, for Gomoku, a pass action is not used. For Connect4, since players can only select from $W$ actions, AlphaViT uses only $p(m,0)$.
To decide on a move, AlphaViT employs Monte Carlo Tree Search (MCTS) with Upper Confidence Bound (UCT), using the value and move probabilities computed by the DNN. The MCTS algorithm used in AlphaViT is identical to that of AlphaZero, as detailed in Appendix 1. The hyperparameters for AlphaViT are listed in Appendix 3.
AlphaViD and AlphaVDA
AlphaViT’s DNN has a significant drawback: the size of the policy vector is fixed by the input size of the transformer encoder layer. To address this issue, we propose two improved DNN architectures for AlphaViD and AlphaVDA. Although their DNNs share many similarities with AlphaViT’s DNN, they employ both a transformer encoder and a transformer decoder to calculate the value and move probability.
In AlphaViD and AlphaVDA, the input board is linearly embedded using a convolutional layer and fed into a transformer encoder, similar to AlphaViT. The input embedding sequence is defined as follows:
${\mathbf{z}}_{0}=[\bm{x}_{\mathrm{value}};\bm{x}_{\mathrm{game}};\bm{x}^{1}_{p}\bm{\mathrm{E}};...;\bm{x}^{N_{p}}_{p}\bm{\mathrm{E}}],$ (6)
where the sequence consists of a value token, a game token, and patch embeddings. The embedding sequence in AlphaViD and AlphaVDA differs from that in AlphaViT. The estimated value is obtained from the value head that processes the output embedding corresponding to the value embedding from the last layer of the transformer encoder. In DNNs of AlphaViD and AlphaVDA, the encoder output is only used directly for value estimation, indicating no need for the embedding size to match the action-space size.
The architecture of AlphaViD’s DNN is shown at the top of Fig. 3. The DNN estimates the move probability using the transformer decoder and MLPp. The input embeddings for the transformer decoder are derived from the outputs of the transformer encoder corresponding to the patch embeddings, which are further processed through a fully connected layer as follows:
${\mathbf{y}}^{\prime}_{0}=\mathrm{MLP}([\bm{z}^{2}_{L};...;\bm{z}^{N_{e}+1}_{L}]),{\mathbf{y}}^{\prime}_{0}\in R^{N_{d}\times D_{d}},$ (7)
where $N_{d}$ is the number of embeddings, and $D_{d}$ is the embedding size. Since the embedding sequence size must match the input size of the transformer decoder, ${\mathbf{y}}^{\prime}_{0}$ is interpolated to ${\mathbf{y}}_{0}\in\mathbb{R}^{N_{a}\times D}$, where $N_{a}$ is the action space size. This interpolation provides flexibility to adjust the action space size depending on the game type and board size. Similar to the original transformer, the transformer decoder receives ${\mathbf{y}}_{0}$ and the output of the transformer encoder ${\mathbf{z}}_{L}$. Finally, MLPp calculates the move probabilities from the output of the last layer of the transformer decoder ${\mathbf{y}}_{L}$. The probability of action $a_{n}$ is given by:
$p(a_{n})=\mathrm{Sigmoid}(\mathrm{MLP}_{p}(\mathrm{LN}({\mathbf{y}}_{L})))(n),$ (8)
where $a_{n}$ is the $n$th action.
The architecture of AlphaVDA’s DNN is shown at the bottom of Fig. 3. The DNN has a similar architecture to AlphaViD’s DNN but uses learnable embeddings ${\mathbf{y}}^{\prime}_{0}$ as input to the transformer decoder. While the sizes of these learnable embeddings are fixed, they are interpolated to ${\mathbf{y}}_{0}$ to match the action space size, ensuring compatibility and adaptability across different game configurations.
To decide on a move, AlphaViD and AlphaVDA use Monte Carlo Tree Search (MCTS) with Upper Confidence Bound (UCT), using the value and move probabilities predicted by their DNNs. The MCTS algorithm implemented in these agents is identical to that of AlphaZero, as detailed in Appendix 1. The hyperparameters for AlphaViD and AlphaVDA are listed in Appendix 3.
4 Experimental setup
4.1 Games
This study evaluates the performance of AlphaViT and AlphaViD across six games: Connect4, Connect4 5x4, Gomoku, Gomoku 6x6, Othello, and Othello 6x6. These games are two-player, deterministic, zero-sum games with perfect information. Connect4, published by Milton, is a connection game played on a $7\times 6$ board. The players take turns dropping discs onto the board. A player wins by forming a straight line of four discs horizontally, vertically, or diagonally. Connect4 5x4 is a variant of Connect4 with a $5\times 4$ board. Gomoku is a connection game in which players place stones on a board to form a line of five stones in a row, either horizontally, vertically, or diagonally. This study uses a $9\times 9$ board for Gomoku and a $6\times 6$ board for Gomoku 6x6. Othello (Reversi) is a two-player strategy game played on an $8\times 8$ board. In Othello, the disc is white on one side and black on the other. Players take turns placing a disc with their assigned color facing up. During a game, discs of the opponent’s color are flipped to the current player’s color if they are in a straight line and bounded by the disc just placed and another disc of the current player’s color. Othello 6x6 is played on a $6\times 6$ board in this study.
4.2 Opponents
This study evaluates the performance of AlphaViT, AlphaViD, and AlphaVDA using five different AI methods: AlphaZero, two variants of Monte Carlo Tree Search (MCTS) labeled MCTS100 and MCTS400, Minimax, and Random. AlphaZero is trained using the method described in Appendix 1. The MCTS methods (MCTS100 and MCTS400) were implemented using different numbers of simulations (100 and 400, respectively). The details of MCTSs are provided in Appendix 4. In these MCTS methods, the child nodes are expanded at the fifth visit to a node. Minimax selects a move using the minimax algorithm based on the evaluation table described in Appendix 5. Random selects moves uniformly at random from valid moves.
4.3 Software
AlphaViT, AlphaViD, AlphaVDA, the opponents, and the board games were implemented in Python, using NumPy for linear algebra operations and PyTorch for deep neural networks. The source code is available on GitHub at https://github.com/KazuhisaFujita/AlphaViT for reproducibility and further extension of this work.
5 Results
The results section presents the performance and characteristics of the proposed AI agents: AlphaViT, AlphaViD, and AlphaVDA. These agents are trained on different games (Connect4, Gomoku, and Othello) with two board sizes (small and large). The architectures of these agents vary primarily according to the number of encoder layers, which influences their capability to learn from the game environment. Table 1 shows the number of parameters for each agent configuration. AlphaViT, AlphaViD, and AlphaVDA are tested with different numbers of encoder layers, denoted as ‘L’ followed by a number (e.g., L1, L4, L5, L8). The number of parameters ranges from 11.2 million to 19.9 million, increasing with the number of encoder layers. For comparison, the AlphaZero agent, which serves as the baseline, has 7.1 million parameters.
Each AI agent is trained on specific games with different board sizes. Table 2 categorizes the agents based on the games on which they are trained and the board sizes used during training. The first group includes agents trained on a single game with a large board, denoted as LB. The second group consists of agents trained on a single game with a small board, denoted as SB. The third group comprises agents trained on multiple games, including Connect4, Gomoku, and Othello, with large boards, denoted as Multi. The agents in the third group are trained simultaneously on the three games and can play these three games using a single DNN. In other words, they do not specialize in a specific game. This diversity in training regimes allows us to evaluate the adaptability and generalization capabilities of the agents across different game domains.
Table 1: Encoder Layer Variations and Parameter Sizes in AI Agents
AI agent num of encoder layers num of parameters
AlphaViT L4 4 11.2M
AlphaViD L1 1 11.5M
AlphaVDA L1 1 11.3M
AlphaViT L8 8 19.6M
AlphaViD L5 5 19.9M
AlphaVDA L5 5 19.8M
AlphaZero - 7.1M
Table 2: Board Size and Game Variations in AI Agent Training
AI agents Game Board size
AlphaViT LB, AlphaViD LB, AlphaVDA LB one specific game Large
AlphaViT SB, AlphaViD SB, AlphaVDA SB one specific game Small
AlphaViT Multi, AlphaViD Multi, AlphaVDA Multi Connect4, Gomoku, Othello Large
5.1 Baseline Elo ratings
Table 3 lists the Elo ratings of various AI agents for different games and board sizes. Elo ratings provide a standard measure of relative performance in two-player games and enable systematic comparison of the performance of each AI agent. The evaluated agents include variations of the proposed methods trained on a single game with a large board (LB), a single game with a small board (SB), and multiple games with large boards (Multi). These are compared with other AI agents, including AlphaZero, Monte Carlo Tree Search (MCTS) with different numbers of simulations, Minimax, and a Random agent. The proposed agents and AlphaZero are trained over 1000 iterations. The Elo ratings of all agents are initialized to 1500 and calculated through 50 round-robin tournaments between the agents. The details of the Elo rating calculation are provided in Appendix 6.
AlphaViT L4 LB achieves strong performance, achieving Elo ratings of 1824 in Connect4, 1835 in Gomoku, and 2125 in Othello. However, for Connect4 and Gomoku, AlphaZero outperforms AlphaViT. Interestingly, despite not being trained on small-board games, AlphaViT L4 LB performs as well as or better than Minimax and MCTS100 in Connect4 5x4, Gomoku 6x6, and Othello 6x6. Furthermore, AlphaViT L4 LB exhibits competitive performance for Othello with AlphaViD L1 SB, AlphaVDA L1 SB, and AlphaZero. This suggests that AlphaViT may efficiently utilize the knowledge obtained from training on large-board games.
AlphaViT L4 SB shows Elo ratings of 1732 in Connect4 5x4, 1764 in Gomoku 6x6, and 1969 in Othello 6x6. In these games, AlphaViT L4 SB performs as well as or better than AlphaZero. For Gomoku and Othello, it outperforms Minimax and MCTS even though it is not trained on large board games. This result suggests that AlphaViT L4 SB may effectively utilize knowledge from training on small-board games for large-board games.
AlphaViT L4 Multi, trained simultaneously on Connect4, Gomoku, and Othello, shows strong results in these games, achieving Elo ratings of 2003 in Connect4, 1969 in Gomoku, and 1939 in Othello. Although it is weaker than AlphaZero in Connect4 and Gomoku, AlphaViT L4 Multi outperforms AlphaViT L4 LB (trained on Gomoku only) in Gomoku. Furthermore, AlphaViT L4 Multi exhibits comparable performance to MCTS400 in Gomoku 6x6.
AlphaViD L1 LB performs slightly worse than AlphaViT in the games with the large boards. For Gomoku 6x6 and Othello 6x6, AlphaViD L1 LB outperforms Minimax and MCTS100, suggesting that it effectively transfers knowledge gained from training on large boards. AlphaViD L1 SB achieves Elo ratings of 1742 in Connect4 5x4, 1785 in Gomoku 6x6, and 1852 in Othello 6x6, with performance comparable to AlphaZero and AlphaViT. AlphaViD L1 Multi shows performance close to AlphaViT in Connect4 but is weaker in Gomoku and Othello.
AlphaVDA L1 LB performs worse than AlphaViT for the games with large boards. AlphaVDA L1 SB shows Elo ratings of 1770 in Connect4 5x4, 1871 in Gomoku 6x6, and 1833 in Othello 6x6, with performance comparable to AlphaZero, AlphaViT, and AlphaViD. AlphaVDA L1 Multi is on par with or slightly weaker than AlphaViD L1 Multi.
The Elo ratings presented in Tab. 3 serve as a baseline for the subsequent experiments described in the following sections.
Table 3: Elo Ratings of AI Agents Across Different Games and Board Sizes
Game Connect4 Connect4 5x4 Gomoku Gomoku 6x6 Othello Othello 6x6
AlphaViT L4 LB 1824 1477 1835 1572 2125 1820
AlphaViD L1 LB 1755 1472 1724 1532 1856 1581
AlphaVDA L1 LB 1757 1506 1594 1244 1745 1336
AlphaViT L4 SB 1169 1732 1564 1764 1487 1969
AlphaViD L1 SB 1187 1742 1503 1785 1128 1852
AlphaVDA L1 SB 1213 1770 973.5 1785 1131 1833
AlphaViT L4 Multi 2003 1374 1969 1654 1939 1136
AlphaViD L1 Multi 1743 1295 1538 1392 1695 1379
AlphaVDA L1 Multi 1720 1321 1524 1194 1565 953.2
AlphaZero 2058 1767 2279 1769 1889 1850
Minimax 1033 1337 1484 1403 1127 1385
MCTS100 1260 1512 1153 1414 1192 1372
MCTS400 1550 1581 1169 1675 1365 1642
Random 725.8 1057 697.0 838.6 741.5 883.2
5.2 Variation of Elo rating over training iteration
Figure 4 shows the Elo ratings of the AlphaViT variants (AlphaViT, AlphaViD, AlphaVDA) and AlphaZero over multiple iterations for large and small board configurations in three games: Connect4, Gomoku, and Othello. The left column shows the results for the large board configurations (standard sizes for Connect4, Gomoku, and Othello). Here, the Elo ratings for single-game-trained agents (AlphaViT L4 LB, AlphaViD L1 LB, AlphaVDA L1 LB), multi-game-trained agents (AlphaViT L4 Multi, AlphaViD L1 Multi, AlphaVDA L1 Multi), and AlphaZero were evaluated at iterations 1, 2, 4, 6, 8, 10, 20, 40, 60, 80, 100, 200, 300, 400, 500, 600, 700, 800, 900, 1000, 1100, 1200, 1300, 1400, 1500, 1600, 1700, 1800, 1900, 2000, 2100, 2200, 2300, 2400, 2500, 2600, 2700, 2800, 2900, and 3000. The right column displays the performance on small board configurations (Connect4 5x4, Gomoku 6x6, and Othello 6x6), with evaluations of single-game-trained models (AlphaViT L4 SB, AlphaViD L1 SB, AlphaVDA L1 SB) and AlphaZero at iterations 1, 2, 4, 6, 8, 10, 20, 40, 60, 80, 100, 200, 300, 400, 500, 600, 700, 800, 900, and 1000.
For games with large boards, all single-game-trained models (AlphaViT L4 LB, AlphaViD L1 LB, AlphaVDA L1 LB) show rapid improvements in Elo ratings, reaching high levels around 300 iterations for Othello and 500 iterations for Gomoku. For Connect4, the Elo ratings continue to improve steadily, with gains even after 1000 iterations. The multi-game-trained models (AlphaViT L4 Multi, AlphaViD L1 Multi, and AlphaVDA L1 Multi) exhibit a similar trend in their Elo ratings to the single-game-trained models. Their performance across the different games shows the flexibility of these agents. After the initial rapid improvement phase, the agents show only minor gains, and their Elo ratings fluctuate.
For small boards, AlphaViT L4 SB, AlphaViD L1 SB, and AlphaVDA L1 SB show very rapid improvements in Elo ratings. After peaking, they reach stable Elo ratings for Connect4 5x4 and Gomoku 6x6. However, for Othello 6x6, the Elo ratings exhibit more fluctuations even after reaching high-performance levels. Because the Elo ratings of the agents rapidly increase and become stable, the evaluation iterations are set to be fewer than those used for larger boards.
These results show that AlphaViT, AlphaViD, and AlphaVDA perform well on small boards, achieving high and stable Elo ratings. However, for games with large boards, agents’ Elo ratings exhibit larger fluctuations. This suggests that further optimizations, such as increasing the depth of the DNNs or adjusting hyperparameters, may be necessary to enhance game-playing skills on larger boards.
Figure 4: Elo rating progression over training iterations for AlphaViT, AlphaViD, and AlphaVDA in large- and small-board games. The left column shows results for large board configurations (Connect4, Gomoku, and Othello), while the right column displays results for their small board configurations (Connect4 5x4, Gomoku 6x6, and Othello 6x6). Solid lines represent single-game-trained agents, whereas dashed lines indicate multi-game-trained agents.
5.3 Effect of transformer encoder depth on performance
In this subsection, we evaluate the performance of agents with deep encoders: AlphaViT L8, AlphaViD L5, and AlphaVDA L5. Figure 5 shows the variations in Elo ratings for these agents over the iterations. The opponents and evaluation conditions are consistent with those used in previous experiments. For the large board configurations (Connect4, Gomoku, and Othello), the Elo ratings stabilize at about 1000-2000 iterations. In contrast, the Elo ratings stabilize at around 100 iterations for Connect4 5x4 and 200 iterations for Gomoku 6x6. For Othello 6x6, the Elo ratings quickly reach a high level but exhibit some fluctuations.
Table 4 provides the mean Elo ratings for each agent, calculated from 2100 to 3000 iterations for the large board configurations. Figure 6 displays the ratio of the mean Elo rating of the agents with deep encoders to that with shallow encoders. The results demonstrate that deeper architectures generally enhance the performance across all games. Notably, AlphaViD Multi and AlphaVDA Multi show significant improvements, particularly for Gomoku, which reach ratios of 1.220 and 1.165, respectively. These agents also improve the performance for Connect4, with ratios of 1.130 and 1.143, and for Othello, with ratios of 1.105 and 1.146. The single-game-trained AlphaViD and AlphaVDA exhibit significant gains for Connect4 and Othello, while their improvements for Gomoku are more modest. In contrast, the single- and multi-game-trained AlphaViTs show weaker improvements than the other agents. These results suggest that increasing the depth of transformer encoder layers enhances the agent’s performance across different board configurations and game types, with a particularly positive impact on AlphaViD and AlphaVDA.
Figure 5: Elo rating progression over training iterations for AlphaViT, AlphaViD, and AlphaVDA with deep encoder in large- and small-board games. The left column shows results for large board configurations (Connect4, Gomoku, and Othello), while the right column displays results for their small board configurations (Connect4 5x4, Gomoku 6x6, and Othello 6x6). Solid lines represent single-game-trained agents, whereas dashed lines indicate multi-game-trained agents.
Table 4: Mean Elo Ratings from 2100 to 3000 iterations
Game Connect4 Gomoku Othello
AlphaViT L4 LB 1966 1753 2007
AlphaViT L8 LB 2028 1862 2046
AlphaViD L1 LB 1814 1656 1756
AlphaViD L5 LB 1902 1704 1928
AlphaVDM L1 LB 1802 1570 1807
AlphaVDM L5 LB 1996 1602 1992
AlphaViT L4 Multi 1897 1917 1964
AlphaViT L8 Multi 2015 2037 1965
AlphaViD L1 Multi 1690 1611 1711
AlphaViD L5 Multi 1909 1966 1891
AlphaVDA L1 Multi 1721 1583 1686
AlphaVDA L5 Multi 1968 1844 1933
AlphaZero 2068 2187 1890
Figure 6: The ratio of Elo ratings between deep and shallow DNNs for AlphaViT, AlphaViD, and AlphaVDM across Connect4, Gomoku, and Othello. Ratios greater than 1.0 indicate superior performance of agents with deep DNNs compared to those with shallow DNNs.
5.4 Effect of fine-tuning from small board games
Figure 7 illustrates the Elo ratings of AlphaViT L8, AlphaViD L5, and AlphaVDA L5 with fine-tuned DNNs and those without fine-tuning (initialized with random weights) across three games: Connect4, Gomoku, and Othello. The fine-tuned DNNs are initialized using weights from the single-game-trained DNNs trained on the small-board configuration at 200 iterations. For example, the fine-tuned DNN for Connect4 is initialized with weights from the DNN trained on Connect4 5x4. In contrast, agents with non-fine-tuned DNNs are initialized with random weights.
The results demonstrate that the agents with the fine-tuned DNNs consistently achieve higher Elo ratings than those with randomly initialized DNNs, particularly for Gomoku and Othello. The performance difference is pronounced in the early training iterations, with fine-tuned DNNs exhibiting rapid improvements and stabilization for Gomoku and Othello. For Gomoku, the advantage of fine-tuning is obvious, as the agents with the fine-tuned DNNs consistently outperform those with the non-fine-tuned DNNs throughout the training process. Some signs of overfitting appear in AlphaViD L5 and AlphaVDA L5 for Gomoku, as their Elo ratings gradually decline from approximately 1000 to 2000 iterations. However, their performance improves after 2000 iterations. In Connect4, the impact of fine-tuning is less significant, with no substantial difference in Elo ratings between agents using fine-tuned and non-fine-tuned DNNs.
Figure 7: Elo rating progression over training iterations of the agents (AlphaViT L8, AlphaViD L5, and AlphaVDA L5) with finetuned and randomly initialized DNNs for three board games (Connect4, Gomoku, and Othello). The three columns show the results for AlphaViT L8, AlphaViD L5, and AlphaVDA L5, respectively. The three rows show the results for Connect4, Gomoku, and Othello.
6 Conclusion and discussion
In this paper, we present AlphaViT, AlphaViD, and AlphaVDA, which are novel game-playing AI agents designed to overcome the limitations of AlphaZero using Vision Transformers (ViT). Unlike AlphaZero, which is restricted to fixed board sizes, our proposed agents demonstrate adaptability, effectively handle variations in board size, and exhibit flexibility across games. Furthermore, these agents can simultaneously train on and play multiple games, such as Connect4, Gomoku, and Othello, within a single shared neural network. The performance of these multitask agents surpasses that of traditional game AI algorithms and, in some cases, approaches the performance of AlphaZero.
The results of this study show that AlphaViT, AlphaViD, and AlphaVDA outperform traditional methods such as Minimax and MCTS across all tested scenarios. Although AlphaZero remains the top performer in some cases, particularly for games with larger boards, the proposed agents exhibit competitive performance. In Othello, AlphaViT, AlphaViD with a deep DNN (AlphaViD L5), and AlphaVDA with a deep DNN (AlphaVDA L5) achieve Elo ratings equal to or higher than AlphaZero. Furthermore, multigame-trained variants perform on par with or better than single-game-trained variants.
AlphaViT, AlphaViD, and AlphaVDA show strong adaptability across different games and board sizes. The agents with DNNs trained on a single game often achieve performance comparable to traditional game algorithms, such as Minimax and MCTS, even when playing on board sizes on which they are not trained. Moreover, multi-game-trained agents frequently perform on par with or surpass their single-game-trained counterparts. The agents with fine-tuned DNNs trained on small-board games consistently achieve performance equal to or better than that of agents with non-fine-tuned DNNs. In the case of Gomoku, using pre-trained weights from small-board games significantly accelerates convergence and enhances the final performance. This suggests effective knowledge transfer between different board sizes, mirroring human learning processes which skills acquired in simpler game variants (e.g., 9x9 Go) can be applied to more complex versions (e.g., 19x19 Go). Such adaptability suggests that the proposed agents may have significant potential for advancing the development of multitask AI.
Comparing the three proposed agents reveals that AlphaViT L4 outperforms both AlphaViD L1 and AlphaVDA L1, despite having an equivalent number of parameters. This difference in performance may be attributed to the fewer transformer encoder layers in AlphaViD and AlphaVDA. This observation is further supported by the fact that AlphaViD L5 and AlphaVDA L5 exhibit performance comparable to AlphaViT L4. The simpler architecture of AlphaViT, consisting solely of encoder layers, may lead to more efficient performance in certain games, even when it has fewer parameters than its variants. However, this simplicity constrains AlphaViT’s flexibility, as its output size is fixed to the number of input embeddings, limiting its applicability to games beyond the classic board games. In contrast, including a decoder layer in AlphaViD and AlphaVDA allows for dynamic adjustment of the policy vector size, providing greater adaptability to games with varying action spaces. This architectural flexibility makes AlphaViD and AlphaVDA versatile candidates for handling more complex games or environments with continuous action spaces.
Our results also relate to those of Soemers et al. (2023), who highlighted the importance of transferring trained policies and value functions across games with varying board sizes and action spaces. Similarly, our agents, equipped with DNNs that utilize weights fine-tuned from a small-board game or trained simultaneously on multiple games, demonstrate enhanced gameplay skills. This suggests that knowledge obtained from small-board games and other games is used efficiently. The ability to train on multiple games concurrently may also help these agents avoid overfitting, further enhancing their generalization capabilities.
Although our proposed agents with transformer-based DNNs are slightly weaker than AlphaZero, we believe they have the potential for further improvement through enhanced training techniques. For example, an efficient decay of the learning rate during training iterations (Silver et al., 2016, 2016, 2018) and warm-start method (Wang et al., 2020, 2021) will improve performance. Increasing the number of self-plays per iteration may also be effective.
In future work, we plan to extend these architectures to a broader range of games, including those with more complex rules and stochastic elements. In addition, we aim to incorporate the flexibility of ViT-based DNNs into other deep reinforcement learning frameworks, such as deep Q-networks, to create AI agents capable of playing a wider variety of games, including video games, with enhanced adaptability.
Appendix
1 AlphaZero
AlphaZero integrates a deep neural network (DNN) with Monte Carlo Tree Search (MCTS), as illustrated in Fig. 1. The DNN processes input representing the current board state and the player, producing an estimated state value and move probabilities. MCTS then uses these outputs to identify the optimal move. Similarly, AlphaViT, AlphaViD, and AlphaVDA adopt this foundational framework.
1.1 Deep neural network in AlphaZero
AlphaZero’s deep neural network (DNN) predicts the value $v(s)$ and the move probability $\bm{p}$ with components $p(a|s)$ for each action $a$, given a state $s$. In the board game context, $s$ and $a$ represent the board state and the move, respectively. In the context of board games, $s$ represents the board state, and $a$ denotes a move. The DNN receives input representing the current board state and the current player’s disc color. Fig. A1 illustrates the DNN architecture, which consists of a Body (residual blocks) and a Head (value and policy heads). The value head outputs the estimated state value $v(s)$, while the policy head produces the move probabilities $\bm{p}$.
The input to the DNN is an $H\times W\times(2T+1)$ image stack that contains $12T+$ 1 binary feature planes of size $H\times W$. Here, $H\times W$ refers to the board size, and $T$ is the number of histories. The first $T$ feature planes represent the occupancy of the player’s discs, with a feature value of 1 indicating that a disc occupies the corresponding cell, and 0 otherwise. Similarly, the following $T$ feature planes represent the occupancy of the opponent’s discs. The last feature plane represents the disc color of the current player, with the disc colors of the first and second players represented by 1 and -1, respectively.
Figure A1: Deep neural network of AlphaZero.
1.2 Monte Carlo tree search in AlphaZero
This subsection provides an explanation of the Monte Carlo Tree Search (MCTS) algorithm used in AlphaZero. Each node in the game tree represents a game state, and each edge $(s,a)$ represents a valid action from that state. The edges store a set of statistics: $\{N(s,a),W(s,a),Q(s,a),p(s,a)\}$, where $N(s,a)$ is the visit count, $W(s,a)$ is the cumulative value, $Q(s,a)=W(s,a)/N(s,a)$ is the mean value, and $p(s,a)$ is the move probability.
The MCTS for AlphaZero consists of four steps: Select, Expand and Evaluate, Backup, and Play. A simulation is defined as a sequence of Select, Expand and Evaluate, and Backup steps, repeated $N_{\mathrm{sim}}$ times. Play is executed after $N_{\mathrm{sim}}$ simulations.
In the Select, the tree is searched from the root node $s_{\mathrm{root}}$ to the leaf node $s_{L}$ at time step $L$ using a variant of the PUCT algorithm. At each time step $t<L$, the selected action $a_{t}$ has the maximum score, as described by the following equation:
$a_{t}=\mathop{\rm arg~{}max}\limits_{a}(Q(s_{t},a)+C_{\mathrm{puct}}P(s_{t},a)\frac{\sqrt{N(s_{t})}}{1+N(s_{t},a)}),$ (9)
where $N(s_{t})$ is the number of parent visits and $C_{\mathrm{puct}}$ is the exploration rate. In this study, $C_{\mathrm{puct}}$ is constant, whereas in the original AlphaZero, $C_{\mathrm{puct}}$ increases slowly with search time. Additionally, when the parent node is the root node, the node selection is performed using an $\epsilon$-greedy algorithm based on the UCT scores.
In the Expand and Evaluate, the DNN evaluates the leaf node and outputs $v(s_{l})$ and $\bm{p}_{a}(s_{l})$. If the leaf node is a terminal node, $v(s_{L})$ is the color of the winning player’s disc. The leaf node is expanded and each edge $(s_{L},a)$ is initialized to $\{N(s_{L},a)=0,W(s_{L},a)=0,Q(s_{L},a)=0,P(s_{L},a)=p_{a}\}$.
In the Backup, the visit counts and values are updated in a backward pass through each step, $t\leq L$. The visit count is incremented by 1, $N(s_{t},a_{t})\leftarrow N(s_{t},a_{t})+1$, and the cumulative and average values are updated, $W(s_{t},a_{t})\leftarrow W(s_{t},a_{t})+v$, $Q(s_{t},a)\leftarrow W(s_{t},a_{t})/N(s_{t},a_{t})$.
Finally, in the Play, AlphaZero selects the action corresponding to the most visited edge from the root node.
2 Training
AlphaViT, AlphaViD, AlphaVDA, and AlphaZero use the same training scheme. The training process consists of three main components: Self-play, Augmentation, and Update, which are iterated $N_{\mathrm{iter}}$ times. This training algorithm is a modified version of the original AlphaZero, adapted to allow training on a single computer.
During the Self-play phase, an agent plays against itself $N_{\mathrm{self}}$ times. For the first $T_{\mathrm{opening}}$ turns, actions are stochastically selected from valid moves based on the softmax policy as follows:
$p(a\mid s)=\exp(N(s,a)/\tau)/\Sigma_{b}\exp(N(s,b)/\tau),$ (10)
where $\tau$ is a temperature parameter that controls the exploration. This stochastic exploration enables the agent to explore new and potentially better actions. After $T_{\mathrm{opening}}$, the most visited action is selected. Through Self-play, we collect board states, winners, and search probabilities. The search probabilities represent the probabilities of selecting valid moves at the root node in MCTS.
In the Augmentation phase, the dataset derived from Self-play is augmented by introducing symmetries specific to the game variant (e.g., two symmetries for Connect4 and eight for Othello and Gomoku). This augmented data is added to a queue with a capacity of $N_{\mathrm{queue}}$ states to form the training dataset.
For the first update iteration, the training data queue is filled with data generated by self-play using MCTS100, which is then augmented. In subsequent iterations, new data generated by Self-play are added to the training data queue. To simultaneously learn multiple games, we prepare a separate training data queue for each game.
During the Update phase, the deep neural network (DNN) is trained using mini-batch stochastic gradient descent with $N_{\mathrm{batch}}$ batch size and $N_{\mathrm{epochs}}$ epochs. The optimization process includes momentum and weight decay. The loss function $l$ combines the mean squared error between the predicted value $v$ and the winner’s disc color $c_{\mathrm{win}}$, and the cross-entropy loss between the search probabilities $\bm{\pi}$ and the predicted move probabilities $\bm{p}$:
$l=(c_{\mathrm{win}}-v)^{2}-\bm{\pi}^{\mathrm{T}}\log\bm{p}.$ (11)
To train multiple games simultaneously, mini-batches are generated from the respective training data queue of each game. During the Update phase, mini-batches are sampled from these individual queues and used to update the DNN. For example, when an agent simultaneously trains Connect4, Gomoku, and Othello, a mini-batch of Connect4, a mini-batch of Gomoku, and a mini-batch of Othello are used orderly for the update.
3 Parameters
The hyperparameters for AlphaViT, AlphaViD, AlphaVDA, and AlphaZero are detailed in Table A1. The weight decay and momentum values are consistent with those specified in the AlphaZero pseudocode (Silver et al., 2018). All other parameters for AlphaZero are consistent with the previous implementation (Fujita, 2022). The hyperparameters of the other models were carefully hand-tuned to optimize their performance.
Table A2 lists the game-specific hyperparameters for AlphaViT, ALphaViD, AlphaVDA, and AlphaZero. The number of MCTS simulations ($N_{\mathrm{sim}}$) ranges from 200 to 400, depending on the game and board size. The number of self-play games per iteration is set to 30 for Connect4 variants and 10 for Gomoku and Othello variants. The opening phase ($T_{\mathrm{opening}}$) specifies the number of initial moves using softmax decision-making with a temperature parameter ($\tau$) that is adjusted based on the game and board size. For AlphaViT, AlphaViD, and AlphaVDA, we employ the AdamW optimizer in PyTorch, with all parameters set to their default values except for the learning rate. These hyperparameters were also carefully hand-tuned.
Table A1: Hyperparameters of AlphaViT, AlphaViD, AlphaVDA, and AlphaZero
Parameter AlphaViT AlphaViD AlphaVDA AlphaZero
$C_{\mathrm{puct}}$ 1.25 1.25 1.25 1.25
$\epsilon$ 0.2 0.2 0.2 0.2
$T$ 1 1 1 1
$N_{\mathrm{queue}}$ 100000 100000 100000 100000
$N_{\mathrm{epoch}}$ 1 1 1 1
Optimizer AdamW AdamW AdamW SGD
Batch size 1024 1024 1024 1024
Learning rate 0.0001 0.0001 0.0001 0.05
Momentum - - - 0.9
Weight decay 0.01 0.01 0.01 0.0001
Patch size 5 5 5 -
Stride for making patches 1 1 1 -
Embedding size of encoder 512 512 512 -
Dim feedforward of encoder 1024 1024 1024 -
Num of encoder heads 8 8 8 -
Size of positional embeddings $512\times 256$ $512\times 256$ $512\times 256$
Num of decoder layers - 1 1 -
Embedding size of decoder - 512 512 -
Dim feedforward of decoder - 1024 1024 -
Num of decoder head - 8 8 -
Size of action embeddings - - 256 -
Dropout rate 0.1 0.1 0.1 -
Num of residual blocks - - - 6
Kernel size - - - 3
Num of filters - - - 256
Table A2: Game-specific Hyperparameters for AlphaViT, AlphaViD, AlphaVDA, and AlphaZero
Connect4 Connect4 5x4 Gomoku Gomoku 66 Othello Othello 66
Num of simulations 200 200 400 200 400 200
Num of self-play 30 30 10 10 10 10
$T_{\mathrm{opening}}$ 4 4 8 4 8 4
$\tau$ 100 100 40 20 80 40
4 Monte Carlo tree search
MCTS is a best-first search method that does not require a predefined evaluation function (Winands, 2017). It operates through four strategic steps: Selection step, Playout step, Expansion step, and Backpropagation step (Winands, 2017). In the Selection step, the tree is traversed from the root node to a leaf node. Child node $c$ of parent node $p$ is selected to maximize the following score:
$\mathrm{UCT}=q_{c}/N_{c}+C\sqrt{\frac{\ln(N_{p}+1)}{N_{c}+\varepsilon}},$ (12)
where $q_{c}$ is the cumulative value of $c$, $N_{p}$ is the visit count of $p$, $N_{c}$ is the visit count of $c$, and $\varepsilon$ is a constant value to avoid division by zero. In the Playout step, a valid move is selected at random until the end of the game is reached. In the Expansion step, child nodes are added to the leaf node when the visit count of the leaf node reaches $N_{\mathrm{open}}$. In the Backpropagation step, the result of the playout is propagated along the path from the leaf node to the root node. If the MCTS player itself wins, loses, and draws, $q_{i}\leftarrow q_{i}+1$, $q_{i}\leftarrow q_{i}-1$, and $q_{i}\leftarrow q_{i}$, respectively. $q_{i}$ is the cumulative value of node $i$. The sequence of these four steps constitutes a single simulation. The simulation process is repeated $N_{\mathrm{sim}}$ times. After all simulations, MCTS selects the action corresponding to the most visited child node of the root. In this study, $C=0.5$, $\varepsilon=10^{-7}$, and $N_{\mathrm{open}}=5$.
5 Minimax algorithm
The minimax algorithm is a fundamental game-tree search technique that determines the optimal action by evaluating the best possible outcome for the current player. Each node in the tree contains a state, player, action, and value. The algorithm creates a game tree with a depth of $N_{\mathrm{depth}}=3$. The root node corresponds to the current state and minimax player. Next, the states corresponding to leaf nodes are evaluated. Then, the algorithm propagates the values from the leaf nodes to the root node. If the player corresponding to the node is the opponent, the value of the node is the minimum value of its child nodes. Otherwise, the value of the node is the maximum value of its child nodes. Finally, the algorithm selects the action corresponding to the root’s child node with the maximum value. The evaluation of leaf nodes is tailored to each game.
For the Connect4 variants, the values of the connections of two and three same-colored discs are $R\times c_{\mathrm{disc}}c_{\mathrm{minimax}}$ and $R^{2}\times c_{\mathrm{disc}}c_{\mathrm{minimax}}$, respectively, where $R$ is the base reward, and $c_{\mathrm{disc}}$ and $c_{\mathrm{minimax}}$ are the colors of the connecting discs and the minimax player’s disc, respectively. The value of a node is the sum of the values of all connections on the corresponding board. The terminal nodes have a value of $R^{3}c_{\mathrm{win}}c_{\mathrm{minimax}}$, where $c_{\mathrm{win}}$ is the color of the winner, and $R=100$.
For the Gomoku variants, the values of the connections of two, three, and four same-colored discs are $R\times c_{\mathrm{disc}}c_{\mathrm{minimax}}$, $R^{2}\times c_{\mathrm{disc}}c_{\mathrm{minimax}}$, and $R^{3}\times c_{\mathrm{disc}}c_{\mathrm{minimax}}$, respectively. The value of a node is the sum of the values of all connections on the corresponding board. The terminal nodes have values of $R^{4}c_{\mathrm{win}}c_{\mathrm{minimax}}$ and $R=100$.
For the Othello variants, the value of a node is calculated using the following equation:
$E=\sum_{x}\sum_{y}v(x,y)o(x,y)c_{\mathrm{minimax}},$ (13)
where $v_{(}x,y)$ is the value of cell $(x,y)$ and $o(x,y)$ is the occupancy of cell $(x,y)$. For 6x6 Othello and 8x8 Othello, the minimax algorithm evaluates each cell using Eq. 14 and 15, respectively. $o(x,y)$ is $1$, $-1$, and 00 if cell $(x,y)$ is occupied by the first player’s disc, the second player’s disc, and empty, respectively. For the Othello variants, the algorithm expands the tree to the terminal nodes after the last six turns. The terminal nodes have a value of $E_{\mathrm{end}}=1000c_{\mathrm{win}}c_{\mathrm{minimax}}$.
$v_{6\times 6}=\left(\begin{matrix}30&-5&2&2&-5&30\\ -5&-15&3&3&-15&-5\\ 2&3&0&0&3&2\\ 2&3&0&0&3&2\\ -5&-15&3&3&-15&-5\\ 30&-5&2&2&-5&30\\ \end{matrix}\right).$ (14)
$v_{8\times 8}=\left(\begin{matrix}120&-20&20&5&5&20&-20&120\\ -20&-40&-5&-5&-5&-5&-40&-20\\ 20&-5&15&3&3&15&-5&20\\ 5&-5&3&3&3&3&-5&5\\ 5&-5&3&3&3&3&-5&5\\ 20&-5&15&3&3&15&-5&20\\ -20&-40&-5&-5&-5&-5&-40&-20\\ 120&-20&20&5&5&20&-20&120\\ \end{matrix}\right).$ (15)
6 Elo rating
Elo rating is a widely used metric for evaluating the relative performance of players in two-player games. This allows us to estimate the probability of one player defeating another based on their current ratings. Given two players A and B with Elo ratings $e(A)$ and $e(B)$, respectively, the probability that player A will defeat player B, denoted $p(A\mathrm{defeats}B)$, is calculated using the following formula:
$p(A~{}\mathrm{defeats}~{}B)=1/(1+10^{(e(B)-e(A))/400}).$ (16)
After a series of $N_{G}$ games between players A and B, player A’s Elo rating is updated to a new value $e^{\prime}(A)$ based on their performance:
$e^{\prime}(A)=e(A)+K(N_{\mathrm{win}}-N_{G}\times p(A~{}\mathrm{defeats}~{}B)),$ (17)
where $N_{\mathrm{win}}$ is the number of times player A has won, and $K$ is a factor that determines the maximum rating adjustment after a single game. In this study, $K=8$.